# BasicAVSR: Arbitrary-Scale Video Super-Resolution via Image Priors and Enhanced Motion Compensation

### v1

@@ 4.2 Comparison with State-of-the-art Methods
We compare BasicAVSR with state-of-the-art AISR and AVSR methods. For AISR, we choose methods from three categories: 1) learnable adaptive filter-based upsampling, including ArbSR [57] and EQSR [59], 2) implicit neural representation-based upsampling, including LTE [30], CLIT [9], OPE [53], and 3) Gaussian splatting-based upsampling, including GaussianSR [21] and ContinuousSR [45]. For AVSR, we compare with VideoINR [12], MoTIF [11], SAVSR [31], ST-AVSR [49] and BF-STVSR [25]. All competing methods have been finetuned on the REDS dataset for a fair comparison, and we evaluate their generalization ability on Vid4 [36] and further assess robustness to unknown degradations by testing on GoPro [43] with a random degradation pipeline [8] and real-world data. More video results are available at the link Video Results.
@@
ArbSR [57] 27.43/0.798/0.239 16.6 887.3 0.651
EQSR [59] 28.00/0.815/0.228 11.6 1743.2 0.921
RDN [61] LTE [30] 28.02/0.805/0.233 22.5 2011.3 0.519
CLIT [9] 28.02/0.805/0.238 37.7 7341.9 1.655
OPE [53] 27.90/0.798/0.242 22.1 1003.7 0.266
GaussianSR [21] 27.97/0.801/0.240 23.2 1576.4 0.712
ContinuousSR [45] 28.04/0.805/0.236 26.0 1980.1 0.319
SwinIR [32] LTE [30] 28.09/0.806/0.231 12.1 1692.8 0.729
CLIT [9] 28.10/0.806/0.237 27.3 7022.3 1.928
OPE [53] 28.02/0.802/0.240 11.7 684.0 0.438
GaussianSR [21] 28.06/0.804/0.236 12.8 1257.9 0.923
ContinuousSR [45] 28.09/0.807/0.233 15.6 1661.6 0.502
VideoINR [12] 27.89/0.802/0.221 11.3 1676.5 0.676
MoTIF [11] 28.02/0.810/0.219 12.6 2826.2 1.132
BF-STVSR [25] 28.14/0.812/0.213 13.5 1876.4 1.003
SAVSR [31] 29.67/0.849/0.193 11.5 1148.0 0.817
ST-AVSR [49] 29.70/0.852/0.195 27.9 296.8 0.101
@@ 4.2.2 Generalization on Vid4
All models trained on REDS are directly applicable to Vid4, which serves as a generalization test. The quantitative results, listed in Table II, indicate that BasicAVSR surpasses all competing methods by wide margins in terms of PSNR, SSIM and LPIPS across varying scaling factors. A closer look is provided in Fig. 6, illustrating the PSNR and LPIPS variations for different scaling factors. It is evident that MoTIF fails to achieve satisfactory SR performance for non-integer and asymmetric scales. This issue is mainly due to the pixel misalignment between the SR frames and ground-truth frames, leading to oscillating PSNR values. Such oscillation is less pronounced in terms of LPIPS as it offers some degree of robustness to misalignment through the VGG feature hierarchy. As for ST-AVSR, it degrades gracefully with increasing scaling factors, including non-integer and asymmetric ones. Our BasicAVSR significantly boosts the generalization of ST-AVSR, achieving up to 1.8 dB PSNR gains in super-resolution across various scaling factors compared to ST-AVSR.
@@
Qualitative results are shown in Fig. 7, where we find that BasicAVSR consistently produces natural and visually pleasing SR outputs. Particularly in high scale factor scenarios, the improvement is pronounced. As shown in the last row of Fig. 7, our method can accurately reconstruct letters, while ST-AVSR fails to recover fine details in the calendar. The refined method, with better alignment accuracy and the strong temporal modeling of bidirectional RNN, is more adept at reconstructing both non-structured and structured texture. Additionally, Fig. 8 compares temporal consistency by unfolding one column of pixels as indicated by the pink dashed line along the temporal dimension. The temporal profiles of the competing methods appear blurry and zigzagging, indicating temporal flickering artifacts. In contrast, the temporal profile of BasicAVSR is closer to the ground-truth, with a sharper and smoother visual appearance.
@@ 4.2.3 Generalization to Unseen Degradation Models
A practical AVSR method must be effective under various, potentially unseen degradations. To evaluate this, we generate test video sequences by incorporating more complex video degradations [8], such as noise and video compression before bicubic downsampling, which are absent from the training data. We applied the aforementioned pipeline to the test set of GoPro to create a test set with unseen degradations. Taking $\times 4$ super-resolution as an example, the results are shown in Table III. Our method demonstrates superior generalization compared to existing approaches. We also conducted a comprehensive comparison of all methods in terms of parameters, computational cost, and runtime using an NVIDIA RTX A6000 GPU. Our ST-AVSR and BasicAVSR not only excel in performance but are also the most efficient in processing. Our BasicVSR further improves performance without markedly increasing inference time. Fig. 9 presents visual comparison of $\times 4$ SR results. Due to the degradation gap between training and testing, all methods, including BasicAVSR, suffer significantly, resulting in missing details in reconstruction results. Nevertheless, BasicAVSR still produces relatively more natural and less distorted results under unseen degradations, further illustrating the superiority of our method. More real-world video results are available at the link Video Results.

### v2

@@ 4.2 Comparison with State-of-the-art Methods
We compare BasicAVSR with state-of-the-art AISR and AVSR methods. For AISR, we choose methods from three categories: 1) learnable adaptive filter-based upsampling, including ArbSR [wang2021learning] and EQSR [wang2023deep], 2) implicit neural representation-based upsampling, including LTE [lee2022local], CLIT [chen2023cascaded], OPE [song2023ope], and 3) Gaussian splatting-based upsampling, including GaussianSR [hu2025gaussiansr] and ContinuousSR [peng2025pixel]. For AVSR, we compare with VideoINR [chen2022videoinr], MoTIF [chen2023motif], SAVSR [li2024savsr], ST-AVSR [shang2024arbitrary], and BF-STVSR [kim2025bf]. All competing methods have been finetuned on the REDS dataset for a fair comparison, and we evaluate their generalization ability on Vid4 [liu2013bayesian] and further assess robustness to unknown degradations by testing on GoPro [Nah_2017_CVPR] with a random degradation pipeline [chan2022investigating] and real-world data. More video results are available at the link Video Results.
@@
ArbSR [wang2021learning] 27.43/0.798/0.239 16.6 887.3 0.651
EQSR [wang2023deep] 28.00/0.815/0.228 11.6 1743.2 0.921
RDN [zhang2018residual] LTE [lee2022local] 28.02/0.805/0.233 22.5 2011.3 0.519
CLIT [chen2023cascaded] 28.02/0.805/0.238 37.7 7341.9 1.655
OPE [song2023ope] 27.90/0.798/0.242 22.1 1003.7 0.266
GaussianSR [hu2025gaussiansr] 27.97/0.801/0.240 23.2 1576.4 0.712
ContinuousSR [peng2025pixel] 28.04/0.805/0.236 26.0 1980.1 0.319
SwinIR [liang2021swinir] LTE [lee2022local] 28.09/0.806/0.231 12.1 1692.8 0.729
CLIT [chen2023cascaded] 28.10/0.806/0.237 27.3 7022.3 1.928
OPE [song2023ope] 28.02/0.802/0.240 11.7 684.0 0.438
GaussianSR [hu2025gaussiansr] 28.06/0.804/0.236 12.8 1257.9 0.923
ContinuousSR [peng2025pixel] 28.09/0.807/0.233 15.6 1661.6 0.502
VideoINR [chen2022videoinr] 27.89/0.802/0.221 11.3 1676.5 0.676
MoTIF [chen2023motif] 28.02/0.810/0.219 12.6 2826.2 1.132
BF-STVSR [kim2025bf] 28.14/0.812/0.213 13.5 1876.4 1.003
SAVSR [li2024savsr] 29.67/0.849/0.193 11.5 1148.0 0.817
ST-AVSR [shang2024arbitrary] 29.70/0.852/0.195 27.9 296.8 0.101
@@ 4.2.2 Generalization on Vid4
All models trained on REDS are directly applicable to Vid4, which serves as a generalization test. The quantitative results, listed in Table II, indicate that BasicAVSR surpasses all competing methods by wide margins in terms of PSNR, SSIM, and LPIPS across varying scaling factors. A closer look is provided in Fig. 6, illustrating the PSNR and LPIPS variations for different scaling factors. It is evident that MoTIF fails to achieve satisfactory SR performance for non-integer and asymmetric scales. This issue is mainly due to the pixel misalignment between the SR frames and ground-truth frames, leading to oscillating PSNR values. Such oscillation is less pronounced in terms of LPIPS as it offers some degree of robustness to misalignment through the VGG feature hierarchy. As for ST-AVSR, it degrades gracefully with increasing scaling factors, including non-integer and asymmetric ones. Our BasicAVSR significantly boosts the generalization of ST-AVSR, achieving up to 1.8 dB PSNR gains in super-resolution across various scaling factors compared to ST-AVSR.
@@
Qualitative results are shown in Fig. 7, where we find that BasicAVSR consistently produces natural and visually pleasing SR outputs. Particularly in high scale factor scenarios, the improvement is pronounced. As shown in the last row of Fig. 7, our method can accurately reconstruct letters, while ST-AVSR fails to recover fine details in the calendar. The refined method, with better alignment accuracy and the strong temporal modeling of bidirectional RNN, is more adept at reconstructing both non-structured and structured textures. Additionally, Fig. 8 compares temporal consistency by unfolding one column of pixels as indicated by the pink dashed line along the temporal dimension. The temporal profiles of the competing methods appear blurry and zigzagging, indicating temporal flickering artifacts. In contrast, the temporal profile of BasicAVSR is closer to the ground-truth, with a sharper and smoother visual appearance.
@@ 4.2.3 Generalization to Unseen Degradation Models
A practical AVSR method must be effective under various, potentially unseen degradations. To evaluate this, we generate test video sequences by incorporating more complex video degradations [chan2022investigating], such as noise and video compression before bicubic downsampling, which are absent from the training data. We applied the aforementioned pipeline to the test set of GoPro to create a test set with unseen degradations. Taking $\times 4$ super-resolution as an example, the results are shown in Table III. Our method demonstrates superior generalization compared to existing approaches. We also conducted a comprehensive comparison of all methods in terms of parameters, computational cost, and runtime using an NVIDIA RTX A6000 GPU. Our ST-AVSR and BasicAVSR not only excel in performance but are also the most efficient in processing. Our BasicVSR further improves performance without markedly increasing inference time. Fig. 9 presents visual comparison of $\times 4$ SR results. Due to the degradation gap between training and testing, all methods, including BasicAVSR, suffer significantly, resulting in missing details in reconstruction results. Nevertheless, BasicAVSR still produces relatively more natural and less distorted results under unseen degradations, further illustrating the superiority of our method. More real-world video results are available at the link Video Results.